OpenStudy (anonymous):

Medal Solve and Check 5 +3 [ 1 + 2 ( 2x -3 ) ] = 6 ( x+5)

OpenStudy (anonymous):

5+3(1+2(2x−3))=6(x+5) Step 1: Simplify both sides of the equation. 5+3(1+2(2x−3))=6(x+5) Simplify: (Show steps) 12x−10=6x+30 Step 2: Subtract 6x from both sides. 12x−10−6x=6x+30−6x 6x−10=30 Step 3: Add 10 to both sides. 6x−10+10=30+10 6x=40 Step 4: Divide both sides by 6. 6x 6 = 40 6 x= 20 3 Answer: x= 20 3

OpenStudy (anonymous):

so the answer is x = 20/3 ?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

thank you soo much :)

OpenStudy (anonymous):

No problem chica